In a local​ market, the monthly price of Internet access service decreases from ​$35 to ​$25​, and the total quantity of monthly accounts across all Internet access providers increases from 130,000 to 230,000. What is the value price elasticity of​ demand, expressed as a positive​ number? ​(Round your answer to two decimal​ places.)
Answer #1

Given p1=$35 ; p2=$25

Q1=130,000 ; Q2=230,000

%change in quantity=(Q2-Q1)/(Q2+Q1)/2)100

      =(230000-130000)/(230000+130000/2)100

=100000/180000×100

=55.55%

% change in price= (P2-P1)/(P2+P1/2)100

=($25-35/(25+35)/2)100

=-10/30×100

=-33.33%

Price elasticity of demand =%change in quantity/%change in price

=55.55%/-33.33%

=-1.6666666667

Round to two decimals=1.67

Therefore the price of elasticity of demand=1.67